# Toward higher-order infinite time Turing machines: simulational $\Gamma$-machines

### v1

@@ Abstract
Infinite time Turing machines (ITTMs) have been introduced by Hamkins and Lewis in their seminal article [HL00]. The strength of the model comes from a limit rule which allows the ITTM to compute through ordinal stages. This rule is simple to describe and another rule would lead to a different model of ordinal computation. The aim of this article is to define a collection of limit rules for which the models of infinite Turing machine they induce have nice properties, akin to that of the ITTM. Through the analysis of the universal ITTM and of its preponderant role in the study of the ITTM, we devise a set of four constraints. A limit rule satisfying those constraints yields a model of infinite machine for which we can define a universal machine. Adding a fifth constraint, we show that any limit rule definable in set theory which meets those constraints produces a model of infinite machine with the desired properties. Among those, the fact that the supremum of the writable ordinal matches that of the clockable ordinal. That is, with the usual notations, the equality $\lambda=\gamma$ holds for any of those limit rules. Eventually, we provide a counter-example to show that the four first constraints alone are not sufficient.
This article is extracted from the fifth chapter of the thesis [Gir24] of Johan Girardot. It constitutes the first main contribution of this thesis and it aims to lay the groundwork for a generalized study of infinite Turing machines.
@@ 1 Introduction
Infinite time Turing machines (ITTMs) have been first introduced by Hamkins and Lewis in [HL00]. Since then it has been thoroughly studied. Recent surveys can be found, such as [Car19, Wel09], while new studies like [Car20, BDL23] continue to appear. Generalizations can also be found in [Car19, FW11, Koe05]. We reintroduce briefly the model of the ITTM to explain the nature of those different generalizations.
The ITTM is a model of computation that computes through the hierarchy of the ordinals. It is built on the same structure as the classical finite Turing machine. Imagine a finite Turing machine $m$ that somehow managed to compute up to some limit ordinal stage $\alpha$. At this stage, it is in some state $q$, has some real number $x$ (an element of $\left\{0,1\right\}^{\mathbb{N}}$, which we more conveniently write $^{\omega}2$) written on its tape and its head is on some cell $i\in\mathbb{N}$. Following [HL00], we call the tuple of those data the snapshot of the machine at this computation stage. Given this snapshot, the computation step from stage $\alpha$ to $\alpha+1$, that is to say the snaphot at stage $\alpha+1$, is simply determined by the program (or the code) of $m$ using the classical rules of finite Turing machines. And so is the computation up to stage $\alpha+k$ for any finite $k$.
@@ 1.1 Generalizations, state-of-the-art and the place of this work among those
For some machine-wise generalizations, the reader can refer to [Koe05] in which Koepke investigates on enhancements of the basic machine structure: what can be said if the tape is not $\omega$ long but $\alpha$ long for some limit ordinal $\alpha$? This yields $\alpha$-ITTMs and very interestingly provides the missing “mechanical” equivalent of $\alpha$-recursion. Further, it is possible to enhance the size of the tapes to $On$ itself (the class of all ordinals) as done in [KS09] to define “ordinal Turing machines” (OTM). In this case, this yields a model of computations whose behavior and power are closely related to $L$ the Gödel’s constructible universe. For example, Koepke shows how in $L$ every set of ordinals is computable with an OTM from a finite set of ordinal parameters.
The structure of the machine may also be more deeply modified to define infinite time register machine (ITRM). This is done in [KM08]: an ITRM has the structure of a register machine and at a limit stage its current line and its registers values are set to the $\liminf$ of their respective previous values. The register may be of ordinal size or again of size $On$ as described in [KS06]. The monograph [Car19] widely discusses those generalizations.
Even further, the work of Fischbach and Seyfferth in [FS13] could be seen as a radical machine-wise generalization. In this article, they develop an ordinal extension of lambda calculus in which a term can be the $\alpha$-fold composition of another term for $\alpha\in On$. Then, as in previous models of computation, the normal form procedure is an extension of the usual procedure for finite terms and again it relies, at limit stages, on an inferior limit.
As for “rule-wise” generalization, to this date and to the best of our knowledge, the only work published is that of Friedman and Welch in [FW11]. In this article, the authors devise a new limit rule, based on a $\liminf$ working with a dynamically restricted history. More precisely, there is a fourth tape called the “rule tape” and at any limit stage, the value of a cell is set as the inferior limit of a subset of its previous values, which subset is determined by the history of the rule tape. The authors show how this yields a model of computation strictly more powerful than ITTMs but which is equally “well-behaved” and which also exhibits strong links with the constructible hierarchy. We discuss this generalization further in the light of our study of the universal machine in Remark 3.35.
@@
It is clear that not any formula which can be seen as limit rule will give rise to as a well-behaved and interesting model as that of ITTMs. One way to measure the interest of such models is to compare them to the classical ITTM. In the case of ITTM, its power can easily be equated with the largeness of its writable and clockable ordinals and one very good representative of its theorical interest is the equality $\lambda=\gamma$. This equality, whose truth was left open in [HL00] and established in [Wel00a], links two constants of very different nature. Here $\lambda$ is the supremum of the writable ordinals, that is ordinals appearing on the output of some halting ITTM, while $\gamma$ is the supremum of the clockable ordinals, that is ordinals corresponding to the halting time of an ITTM. This equality is fruitful as it provides a link between ITTMs and the constructible hierarchy. For example, with it, we know that all writable sets are written before stage $\gamma=\lambda$ and so are in $L_{\lambda}$ while the fact that all ordinals below $\lambda$ are writable implies that any set in $L_{\lambda}$ is writable.
@@ 1.2 Results and organization of the paper
To begin answering this question, “how and why does the universal ITTM work?”, we start in Section 2 with a brief overview of the ITTM and of common state-of-the-art proof techniques. This will put into light the importante role of simulations and, in particular, of the universal ITTM and provide the main intuitions behind this work. More precisely, we begin this section by proving a slightly new structural equality (which generalizes the equality $\lambda=\gamma$ proved in [Wel00a]) for the usual ITTMs. This proof shows how the universal machine is canonically used in this kind of advanced proofs developed by Welch in [Wel00a]. We then try to understand what is implicitly needed in this proof regarding the universal machine and more generally regarding the possibility of simulating other machines. At the end of this discussion, we exhibit five critical characteristics satisfied by the usual $\limsup$ limit rule. The first four of those form a sufficient condition for a given limit rule to yield a model of ITTM in which a universal machine exists. Again, this section serves as an extended introduction presenting the motivations and intuitions for this article.
@@ 2 Eventually clocking and accidentally clocking
But more interestingly, as done in [HL00], in the infinite setting of ITTMs we can define different degrees of writability for the real numbers. As said, a real number $x$ can be writable if there is a machine that halts (by reaching its distinguished state $q_{halt}$) and, upon halting, which has $x$ written on its output tape. But it can also be eventually writable (sometimes abbreviated e.w.) if there is a machine that never stops and that, at some point, has $x$ written on its output tape and never modifies it afterward. It can be accidentally writable (sometimes abbreviated a.w.) if there is a machine in which $x$ is written on any of its tapes at any stage of its computation, even if it is modified or erased just after. It is easy to see that being writable implies being eventually writable implies being accidentally writable.
@@
and such that the order-type of $\prec_{\alpha}$ is $\alpha$. What motivates this definition is a fundamental result of Hamkins and Lewis, Theorem 2.2 of [HL00], which states that, with this encoding, an ITTM can decide whether a real number encodes a well-order.
So, as ordinals can be written, we may wonder what kind of ordinals can be written, or at least accidentally written. First, given the simple fact that ITTMs have $\omega$ cells, they can only write countable ordinals. Moreover, as there are countably many ITTMs, the writable and the eventually writable ordinals are bounded in $\omega_{1}$; that is their union is countable as well. However, it may not seem so clear whether this fact holds for the accidentally writable ordinals, since a single machine will accidentally write many different ordinals through its computation. Still it happens that those are bounded as well. Hence, as introduced in [HL00] there exists three supremum ordinals, $\Sigma$, the supremum of the accidentally writable ordinals, $\zeta$, the supremum of the eventually writable ordinals and $\lambda$, the supremum of the writable ordinals.
Additionnaly Hamkins and Lewis define the clockable ordinals. An ordinal $\alpha$ is clockable when there is some machine which halts just after carrying out its $\alpha^{th}$ step of computation. That is, through this $\alpha^{th}$ step, it transitions from stage $\alpha$ to stage $\alpha+1$ and, in this latter stage, the machine reached the state $q_{halt}$. Then as previously, there is $\gamma$ the supremum of the clockable ordinals. One important question left open in [HL00] was: does $\lambda=\gamma$? This was answered positively in Theorem 1.1 of [Wel00a] using a universal machine; which now brings us to this machine and its importance.
A universal machine $\mathcal{U}$ is a machine that simulates simultaneously all ITTMs. It was already used by Hamkins and Lewis in Theorem 3.4 of [HL00] and its importance appears even more clearly in [Wel00a]. To illustrate this, we give a proof of a slightly more general equality than that established by Welch, but which relies on the same ideas. As there are three different ways for an ordinal to be written (writable, eventually writable and accidentally writable), we would also can devise three ways for an ordinal to be clockable: and this can be done by saying that: $\alpha$ is eventually clockable (e.c.) when the output of some converging machine $m$ stabilizes at stage $\alpha$, and $\alpha$ is accidentally clockable (a.c.) when there is a computation in which at stage $\alpha$ some real number $x$ appears for the first time in the computation. Then we can write $\eta$ the supremum of the eventually clockable ordinals and $\mathrm{T}$ (capital $\tau$) the supremum of the accidentally clockable ordinals. And with the latter, we can show the following result, with $\Sigma$ being the supremum of the accidentally writable ordinals.
Proposition 2.1 ([Wel00a]).
@@
This proposition, while stated here in a new way, is very close to Corollary 3.5 of [Wel00a] which states that the collection of sets encoded by accidentally writable real numbers is actually $L_{\Sigma}$. In this article from Welch, it was a corollary of the fact that $\lambda=\gamma$ and this fact is itself a corollary of Theorem 2.1 here, as will be proved in a more general setting in Corollary 4.31. Another corollary is the fact that, with previous notations, $\zeta=\eta$. So this downward chain of consequences likely justify this new phrasing. Moreover, as said, the proof of this slightly more general result uses the same technique as that of the main proposition of [Wel00a]. Finally, we take for granted in this proof that there exists a universal ITTM $\mathcal{U}$ and that we can design, in a natural way, machines which simulate other machines. Those fact are actually true and we will come back to those.
@@
First, suppose that $\Sigma>\mathrm{T}$. This implies that $\mathrm{T}$ is accidentally writable by some machine $m_{\mathrm{T}}$. So we consider the following machine $\mathcal{M}$. It will use the code of $\mathrm{T}$ that appears at some point in $m_{\mathrm{T}}$ to write something for the first time after stage $\mathrm{T}$, hence reaching a contradiction. To do this, first, it keeps the first cell of its working tape always set to 0 and the first cell of its output tape always set to $1$. So when using its working tape, it uses cells $i>0$. Then, it simulates $m_{\mathrm{T}}$ on some part of its working tape (again, without using cell $i=0$) and each time an ordinal $\alpha$ is written in the simulation of $m_{\mathrm{T}}$, it counts for at least $\alpha$ steps. Remember that by [HL00, Theorem 2.2] it can recognize real numbers describing well-orders. This is done by trying to count through it, so it takes at least $\alpha$ steps to recognize that a real number encodes a well-order of order type $\alpha$. Then, it copies $\alpha$ on its output tape, without using the first cell that it keeps to $1$. Now, when $\mathrm{T}$ appears for the first time at stage $\alpha_{\mathrm{T}}$: as $\mathrm{T}$ appears for the first time in $m_{\mathrm{T}}$ and $\mathrm{T}$ is the supremum of the a.c., that is of the stages of first appearance in a given machine, we have $\alpha_{\mathrm{T}}<\mathrm{T}$. Then, $\mathcal{M}$ counts for at least $\mathrm{T}$ steps and after those writes $\mathrm{T}$ on its output tape. This is the first time that $\mathrm{T}$ appears on this tape. Also, the whole content of this tape, with the first cell set to $1$ cannot have appeared earlier in the working tape whose first cell is left on 0. Hence the real number it writes was not written before and it is written for the first time at least at stage $\alpha_{\mathrm{T}}+\mathrm{T}\geqslant\mathrm{T}$. And $\alpha_{\mathrm{T}}+\mathrm{T}$ is consequently a.c. which is a contradiction. So $\Sigma\leqslant\mathrm{T}$.
@@
  • Suppose $C_{i}(\Sigma)=0$. Then, by the limsup rule, the cell must have converged to 0 at some least ordinal stage $\alpha<\Sigma$. That is $C_{i}(\alpha)=0$ and stays so up to $\Sigma$. We show that $\alpha<\zeta$ (this is actually Main Proposition of [Wel00a]). We design a machine $\mathcal{M}$ that does the following: it launches a copy $\mathcal{U}_{1}$ of the universal machine and for each ordinal $\beta$ appearing in the computation of $\mathcal{U}_{1}$, that is accidentally written in a machine simulated by $\mathcal{U}_{1}$, $\mathcal{M}$ writes this ordinal on its output tape and simulates $m$ up to this ordinal $\beta$. If $C_{i}^{m}(\beta)=0$, it launches a new copy $\mathcal{U}_{2}$ of the universal machine and each time some ordinal $\beta^{\prime}>\beta$ is appears in $\mathcal{U}_{2}$, it simulates a fresh copy of $m$ up to stage $\beta^{\prime}$ and looks whether the cell $i$ was set back to $1$ between $\beta$ and $\beta^{\prime}$. If it was, $\mathcal{M}$ goes back to its simulation of the universal machine $\mathcal{U}_{1}$ and looks for the next such ordinal $\beta$, that is such that $C_{i}^{m}(\beta)=0$. When $\beta>\alpha$, the machine never find $1$’s anymore in the history of cell $i$ after this $\beta$, as all $\beta^{\prime}>\beta$ generated are strictly between $\alpha$ and $\Sigma$. Hence, $\beta$ is written on the output and $\mathcal{M}$ looks indefinitely for $\beta^{\prime}>\beta$ at which a $1$ appear on cell $i$. And so $\mathcal{M}$ actually eventually writes some $\beta>\alpha$. As such, $\beta$ is eventually writable and we have $\alpha<\beta<\zeta<\Sigma$ and $C_{i}(\zeta)=0$ as well.

### v2

@@ Abstract
Infinite time Turing machines (ITTMs) have been introduced by Hamkins and Lewis in their seminal article [12]. The strength of the model comes from a limit rule which allows the ITTM to compute through ordinal stages. This rule is simple to describe and another rule would lead to a different model of ordinal computation. The aim of this article is to define a collection of limit rules for which the models of infinite Turing machine they induce have nice properties, akin to that of the ITTM. Through the analysis of the universal ITTM and of its preponderant role in the study of the ITTM, we devise a set of four constraints. A limit rule satisfying those constraints yields a model of infinite machine for which we can define a universal machine. Adding a fifth constraint, we show that any limit rule definable in set theory which meets those constraints produces a model of infinite machine with the desired properties. Among those, the fact that the supremum of the writable ordinal matches that of the clockable ordinal. That is, with the usual notations, the equality $\lambda=\gamma$ holds for any of those limit rules. Eventually, we provide a counter-example to show that the four first constraints alone are not sufficient.
This article is extracted from the fifth chapter of the thesis [10] of Johan Girardot. It constitutes the first main contribution of this thesis and it aims to lay the groundwork for a generalized study of infinite Turing machines.
@@ 1 Introduction
Infinite time Turing machines (ITTMs) have been first introduced by Hamkins and Lewis in [12]. Since then it has been thoroughly studied. Recent surveys can be found, such as [1, 3], while new studies like [6, 5] continue to appear. Generalizations can also be found in [1, 9, 13]. We reintroduce briefly the model of the ITTM to explain the nature of those different generalizations.
The ITTM is a model of computation that computes through the hierarchy of the ordinals. It is built on the same structure as the classical finite Turing machine. Imagine a finite Turing machine $m$ that somehow managed to compute up to some limit ordinal stage $\alpha$. At this stage, it is in some state $q$, has some real number $x$ (an element of $\left\{0,1\right\}^{\mathbb{N}}$, which we more conveniently write $^{\omega}2$) written on its tape and its head is on some cell $i\in\mathbb{N}$. Following [12], we call the tuple of those data the snapshot of the machine at this computation stage. Given this snapshot, the computation step from stage $\alpha$ to $\alpha+1$, that is to say the snaphot at stage $\alpha+1$, is simply determined by the program (or the code) of $m$ using the classical rules of finite Turing machines. And so is the computation up to stage $\alpha+k$ for any finite $k$.
@@ 1.1 Generalizations, state-of-the-art and the place of this work among those
For some machine-wise generalizations, the reader can refer to [13] in which Koepke investigates on enhancements of the basic machine structure: what can be said if the tape is not $\omega$ long but $\alpha$ long for some limit ordinal $\alpha$? This yields $\alpha$-ITTMs and very interestingly provides the missing “mechanical” equivalent of $\alpha$-recursion. Further, it is possible to enhance the size of the tapes to $On$ itself (the class of all ordinals) as done in [4] to define “ordinal Turing machines” (OTM). In this case, this yields a model of computations whose behavior and power are closely related to $L$ the Gödel’s constructible universe. For example, Koepke shows how in $L$ every set of ordinals is computable with an OTM from a finite set of ordinal parameters.
The structure of the machine may also be more deeply modified to define infinite time register machine (ITRM). This is done in [14]: an ITRM has the structure of a register machine and at a limit stage its current line and its registers values are set to the $\liminf$ of their respective previous values. The register may be of ordinal size or again of size $On$ as described in [15]. The monograph [1] widely discusses those generalizations.
Even further, the work of Fischbach and Seyfferth in [8] could be seen as a radical machine-wise generalization. In this article, they develop an ordinal extension of lambda calculus in which a term can be the $\alpha$-fold composition of another term for $\alpha\in On$. Then, as in previous models of computation, the normal form procedure is an extension of the usual procedure for finite terms and again it relies, at limit stages, on an inferior limit.
As for “rule-wise” generalization, to this date and to the best of our knowledge, the only work published is that of Friedman and Welch in [9]. In this article, the authors devise a new limit rule, based on a $\liminf$ working with a dynamically restricted history. More precisely, there is a fourth tape called the “rule tape” and at any limit stage, the value of a cell is set as the inferior limit of a subset of its previous values, which subset is determined by the history of the rule tape. The authors show how this yields a model of computation strictly more powerful than ITTMs but which is equally “well-behaved” and which also exhibits strong links with the constructible hierarchy. We discuss this generalization further in the light of our study of the universal machine in Remark 3.35.
@@
It is clear that not any formula which can be seen as limit rule will give rise to as a well-behaved and interesting model as that of ITTMs. One way to measure the interest of such models is to compare them to the classical ITTM. In the case of ITTM, its power can easily be equated with the largeness of its writable and clockable ordinals and one very good representative of its theorical interest is the equality $\lambda=\gamma$. This equality, whose truth was left open in [12] and established in [17], links two constants of very different nature. Here $\lambda$ is the supremum of the writable ordinals, that is ordinals appearing on the output of some halting ITTM, while $\gamma$ is the supremum of the clockable ordinals, that is ordinals corresponding to the halting time of an ITTM. This equality is fruitful as it provides a link between ITTMs and the constructible hierarchy. For example, with it, we know that all writable sets are written before stage $\gamma=\lambda$ and so are in $L_{\lambda}$ while the fact that all ordinals below $\lambda$ are writable implies that any set in $L_{\lambda}$ is writable.
@@ 1.2 Results and organization of the paper
To begin answering this question, “how and why does the universal ITTM work?”, we start in Section 2 with a brief overview of the ITTM and of common state-of-the-art proof techniques. This will put into light the importante role of simulations and, in particular, of the universal ITTM and provide the main intuitions behind this work. More precisely, we begin this section by proving a slightly new structural equality (which generalizes the equality $\lambda=\gamma$ proved in [17]) for the usual ITTMs. This proof shows how the universal machine is canonically used in this kind of advanced proofs developed by Welch in [17]. We then try to understand what is implicitly needed in this proof regarding the universal machine and more generally regarding the possibility of simulating other machines. At the end of this discussion, we exhibit five critical characteristics satisfied by the usual $\limsup$ limit rule. The first four of those form a sufficient condition for a given limit rule to yield a model of ITTM in which a universal machine exists. Again, this section serves as an extended introduction presenting the motivations and intuitions for this article.
@@ 2 Eventually clocking and accidentally clocking
But more interestingly, as done in [12], in the infinite setting of ITTMs we can define different degrees of writability for the real numbers. As said, a real number $x$ can be writable if there is a machine that halts (by reaching its distinguished state $q_{halt}$) and, upon halting, which has $x$ written on its output tape. But it can also be eventually writable (sometimes abbreviated e.w.) if there is a machine that never stops and that, at some point, has $x$ written on its output tape and never modifies it afterward. It can be accidentally writable (sometimes abbreviated a.w.) if there is a machine in which $x$ is written on any of its tapes at any stage of its computation, even if it is modified or erased just after. It is easy to see that being writable implies being eventually writable implies being accidentally writable.
@@
and such that the order-type of $\prec_{\alpha}$ is $\alpha$. What motivates this definition is a fundamental result of Hamkins and Lewis, Theorem 2.2 of [12], which states that, with this encoding, an ITTM can decide whether a real number encodes a well-order.
So, as ordinals can be written, we may wonder what kind of ordinals can be written, or at least accidentally written. First, given the simple fact that ITTMs have $\omega$ cells, they can only write countable ordinals. Moreover, as there are countably many ITTMs, the writable and the eventually writable ordinals are bounded in $\omega_{1}$; that is their union is countable as well. However, it may not seem so clear whether this fact holds for the accidentally writable ordinals, since a single machine will accidentally write many different ordinals through its computation. Still it happens that those are bounded as well. Hence, as introduced in [12] there exists three supremum ordinals, $\Sigma$, the supremum of the accidentally writable ordinals, $\zeta$, the supremum of the eventually writable ordinals and $\lambda$, the supremum of the writable ordinals.
Additionnaly Hamkins and Lewis define the clockable ordinals. An ordinal $\alpha$ is clockable when there is some machine which halts just after carrying out its $\alpha^{th}$ step of computation. That is, through this $\alpha^{th}$ step, it transitions from stage $\alpha$ to stage $\alpha+1$ and, in this latter stage, the machine reached the state $q_{halt}$. Then as previously, there is $\gamma$ the supremum of the clockable ordinals. One important question left open in [12] was: does $\lambda=\gamma$? This was answered positively in Theorem 1.1 of [17] using a universal machine; which now brings us to this machine and its importance.
A universal machine $\mathcal{U}$ is a machine that simulates simultaneously all ITTMs. It was already used by Hamkins and Lewis in Theorem 3.4 of [12] and its importance appears even more clearly in [17]. To illustrate this, we give a proof of a slightly more general equality than that established by Welch, but which relies on the same ideas. As there are three different ways for an ordinal to be written (writable, eventually writable and accidentally writable), we would also can devise three ways for an ordinal to be clockable: and this can be done by saying that: $\alpha$ is eventually clockable (e.c.) when the output of some converging machine $m$ stabilizes at stage $\alpha$, and $\alpha$ is accidentally clockable (a.c.) when there is a computation in which at stage $\alpha$ some real number $x$ appears for the first time in the computation. Then we can write $\eta$ the supremum of the eventually clockable ordinals and $\mathrm{T}$ (capital $\tau$) the supremum of the accidentally clockable ordinals. And with the latter, we can show the following result, with $\Sigma$ being the supremum of the accidentally writable ordinals.
Proposition 2.1 ([17]).
@@
This proposition, while stated here in a new way, is very close to Corollary 3.5 of [17] which states that the collection of sets encoded by accidentally writable real numbers is actually $L_{\Sigma}$. In this article from Welch, it was a corollary of the fact that $\lambda=\gamma$ and this fact is itself a corollary of Theorem 2.1 here, as will be proved in a more general setting in Corollary 4.31. Another corollary is the fact that, with previous notations, $\zeta=\eta$. So this downward chain of consequences likely justify this new phrasing. Moreover, as said, the proof of this slightly more general result uses the same technique as that of the main proposition of [17]. Finally, we take for granted in this proof that there exists a universal ITTM $\mathcal{U}$ and that we can design, in a natural way, machines which simulate other machines. Those fact are actually true and we will come back to those.
@@
First, suppose that $\Sigma>\mathrm{T}$. This implies that $\mathrm{T}$ is accidentally writable by some machine $m_{\mathrm{T}}$. So we consider the following machine $\mathcal{M}$. It will use the code of $\mathrm{T}$ that appears at some point in $m_{\mathrm{T}}$ to write something for the first time after stage $\mathrm{T}$, hence reaching a contradiction. To do this, first, it keeps the first cell of its working tape always set to 0 and the first cell of its output tape always set to $1$. So when using its working tape, it uses cells $i>0$. Then, it simulates $m_{\mathrm{T}}$ on some part of its working tape (again, without using cell $i=0$) and each time an ordinal $\alpha$ is written in the simulation of $m_{\mathrm{T}}$, it counts for at least $\alpha$ steps. Remember that by [12, Theorem 2.2] it can recognize real numbers describing well-orders. This is done by trying to count through it, so it takes at least $\alpha$ steps to recognize that a real number encodes a well-order of order type $\alpha$. Then, it copies $\alpha$ on its output tape, without using the first cell that it keeps to $1$. Now, when $\mathrm{T}$ appears for the first time at stage $\alpha_{\mathrm{T}}$: as $\mathrm{T}$ appears for the first time in $m_{\mathrm{T}}$ and $\mathrm{T}$ is the supremum of the a.c., that is of the stages of first appearance in a given machine, we have $\alpha_{\mathrm{T}}<\mathrm{T}$. Then, $\mathcal{M}$ counts for at least $\mathrm{T}$ steps and after those writes $\mathrm{T}$ on its output tape. This is the first time that $\mathrm{T}$ appears on this tape. Also, the whole content of this tape, with the first cell set to $1$ cannot have appeared earlier in the working tape whose first cell is left on 0. Hence the real number it writes was not written before and it is written for the first time at least at stage $\alpha_{\mathrm{T}}+\mathrm{T}\geqslant\mathrm{T}$. And $\alpha_{\mathrm{T}}+\mathrm{T}$ is consequently a.c. which is a contradiction. So $\Sigma\leqslant\mathrm{T}$.
@@
  • Suppose $C_{i}(\Sigma)=0$. Then, by the limsup rule, the cell must have converged to 0 at some least ordinal stage $\alpha<\Sigma$. That is $C_{i}(\alpha)=0$ and stays so up to $\Sigma$. We show that $\alpha<\zeta$ (this is actually Main Proposition of [17]). We design a machine $\mathcal{M}$ that does the following: it launches a copy $\mathcal{U}_{1}$ of the universal machine and for each ordinal $\beta$ appearing in the computation of $\mathcal{U}_{1}$, that is accidentally written in a machine simulated by $\mathcal{U}_{1}$, $\mathcal{M}$ writes this ordinal on its output tape and simulates $m$ up to this ordinal $\beta$. If $C_{i}^{m}(\beta)=0$, it launches a new copy $\mathcal{U}_{2}$ of the universal machine and each time some ordinal $\beta^{\prime}>\beta$ is appears in $\mathcal{U}_{2}$, it simulates a fresh copy of $m$ up to stage $\beta^{\prime}$ and looks whether the cell $i$ was set back to $1$ between $\beta$ and $\beta^{\prime}$. If it was, $\mathcal{M}$ goes back to its simulation of the universal machine $\mathcal{U}_{1}$ and looks for the next such ordinal $\beta$, that is such that $C_{i}^{m}(\beta)=0$. When $\beta>\alpha$, the machine never find $1$’s anymore in the history of cell $i$ after this $\beta$, as all $\beta^{\prime}>\beta$ generated are strictly between $\alpha$ and $\Sigma$. Hence, $\beta$ is written on the output and $\mathcal{M}$ looks indefinitely for $\beta^{\prime}>\beta$ at which a $1$ appear on cell $i$. And so $\mathcal{M}$ actually eventually writes some $\beta>\alpha$. As such, $\beta$ is eventually writable and we have $\alpha<\beta<\zeta<\Sigma$ and $C_{i}(\zeta)=0$ as well.